OpenStudy (anonymous):

Solve the equation of 2log (2y + 4) = log 9 + 2log (y -1).

OpenStudy (anonymous):

A. 3 B. 5 C. 7 D. 21

OpenStudy (anonymous):

How would i solve this ?

OpenStudy (agent0smith):

Condense the logs on the right side first.

OpenStudy (anonymous):

log9+log2 first or log2(x-1)?

OpenStudy (agent0smith):

first bring the exponents up like this \[\Large \log (2y + 4)^2 = \log 9 + \log (y -1)^2\] then condense logs on the right

OpenStudy (anonymous):

Sorry not sure how to do this could you walk me throught it?

OpenStudy (agent0smith):

Use the log rule log(a) + log(b) = log(a)(b)

OpenStudy (agent0smith):

eg log(x+3) + log(x-2) = log(x+3)(x-2)

OpenStudy (anonymous):

log(9)(2y-1)^2

OpenStudy (agent0smith):

\[\Large \log (2y + 4)^2 = \log 9(y -1)^2 \]now you can just drop the logs on both sides, and solve for y. Logs are one-to-one functions so you can do that here. eg if logx = logy then x = y

OpenStudy (anonymous):

(2y+4)^2=9(2y-1)^2 would be the qeuation to solve

OpenStudy (agent0smith):

Yep